OpenStudy (anonymous):

Let f(x)=x^2+5x, g(x)=6-2x, h(x)=2/x-3; find the following value. f(6)-4= ?

OpenStudy (anonymous):

is it 38??

OpenStudy (anonymous):

Nope :( 66 62 14 or 2

OpenStudy (ash2326):

We have \[\large f(x)=x^2+5x\] We have to find \[\large f(6)-4\] \[f(6)=f(x=6)\] so \[f(6)=6^2+5\times 6=36+30=66\] so \[f(6)-4=66-4=62\]

OpenStudy (anonymous):

Thanks so much ash2326!